OpenStudy (danmac0710):

Please, how can find the square root of: ((3+root(5))/2)

hartnn (hartnn):

it will contain, the root 5 term , right? so let the square root be a+b root 5

hartnn (hartnn):

so we have \(\sqrt{(3+\sqrt5)/2}=a+b\sqrt5\) now square both sides, what u get ?

hartnn (hartnn):

there's another shorter way, of splitting the '3' but you can first try this :)

OpenStudy (jack1):

^^^ what he said ;D

OpenStudy (danmac0710):

thanks guys! That is greaaaaaaaat!

OpenStudy (danmac0710):

hartnn, can you show me the shorter way too please? I did it with the longer method.

hartnn (hartnn):

\(\sqrt{(3+\sqrt5)/2}=(1/2)\sqrt{6+2\sqrt5}=(1/2)\sqrt{1+5+2\sqrt 5}\) note that now,quantity under the square root has form \(a^2+b^2+2ab \\ \quad which =(a+b)^2\) so, \((1/2)\sqrt{1+5+2\sqrt 5}= (1/2)\sqrt{1^2+(\sqrt5)^2+2*1*\sqrt 5} \\ = (1/2)\sqrt{(1+\sqrt5)^2}=(1/2)(1+\sqrt 5)\) got it ?

hartnn (hartnn):

note i split 6 as 1+5

OpenStudy (danmac0710):

Thanks - yes I have it now. Very nice method!